• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 16:00:39 on July23,2003

    Entry number 106341

    keyword=Day Shift Summary

    8:00 Target Replacement Continues

    8:30 Began Septum Training

    9:09 Target Cell Replaced

    11:08 Septum Quench at 408.5 A

    11:30 Patrica and J.P. done with Laser Alignment.

    12:30 Request Beam Permit from MCC.

    12:42 Beam Permit Recieved

    12:55 NPR Measurement on Priapus Cell P=2.9%

    14:15 MCC called we expect pulsed beam soon.

    14:20 We start getting pulsed beam.

    14:50 David L. says that the beam is going through the Compton.

    16:00 Shift over.